### v1

# OpenFACADES: An Open Framework for Architectural Caption and Attribute Data Enrichment via Street View Imagery

Abstract
Building properties, such as height, usage, and material composition, play a crucial role in spatial data infrastructures, supporting applications such as energy simulation, risk assessment, and environmental modeling. Despite their importance, comprehensive and high-quality building attribute data remain scarce in many urban areas. Recent advances have enabled the extraction and tagging of objective building attributes using remote sensing and street-level imagery. However, establishing a method and pipeline that integrates diverse open datasets, acquires holistic building imagery at scale, and infers comprehensive building attributes remains a significant challenge. Among the first, this study bridges the gaps by introducing OpenFACADES, an open framework that leverages multimodal crowdsourced data to enrich building profiles with both objective attributes and semantic descriptors through multimodal large language models. Our methodology proceeds in three major steps. First, we integrate street‐level image metadata from Mapillary with OpenStreetMap geometries via isovist analysis, effectively identifying images that provide suitable vantage points for observing target buildings. Second, we automate the detection of building facades in panoramic imagery and tailor a reprojection approach to convert objects into holistic perspective views that approximate real-world observation. Third, we introduce an innovative approach that harnesses and systematically investigates the capabilities of open-source large vision-language models (VLMs) for multi-attribute prediction and open-vocabulary captioning in building-level analytics, leveraging a globally sourced dataset of 30,180 labeled images from seven cities. Evaluation shows that fine-tuned VLM excel in multi-attribute inference, outperforming single-attribute computer vision models and zero-shot ChatGPT-4o. Further experiments confirm its superior generalization and robustness across culturally distinct regions and varying image conditions. Finally, the model is applied for large-scale building annotation, generating a dataset of 1.2 million images for half a million buildings. This open‐source framework enhances the scope, adaptability, and granularity of building‐level assessments, enabling more fine‐grained and interpretable insights into the built environment. Our dataset and code are available openly at: https://github.com/seshing/OpenFACADES.
keywords:
Building exteriors , Street-level , Volunteered geographic information , ChatGPT , Multi-task learning , SDI
[…]
Abstract
1 Introduction
Buildings, as prominent artifacts within urban settings, serve as vital indicators of the management, transformation, and overall dynamism of the built environment. Their physical characteristics — including geometry, height, function, material, condition, and style — are the key parameters that not only support sustainable urban development but also reflect economic progress and cultural evolution over time (Biljecki et al., 2021). Such rich building-level data has been instrumental in a range of applications, such as urban climate simulations for improved environmental planning (Creutzig et al., 2019), building energy modeling for resource optimization (Kumar et al., 2018; Roth et al., 2020), estimation of urban material stocks for the circular economy (Raghu et al., 2023), and disaster impact assessments to inform effective response and recovery efforts (Westrope et al., 2014). Moreover, these data support more nuanced analyses of population distributions (Schug et al., 2021), socio-economic conditions (Feldmeyer et al., 2020), as well as deeper understanding of the impact on human behaviors (Wang et al., 2016) and public perception (Liang et al., 2024). Hence, more comprehensive and openly accessible geospatial data on building information can enable the formulation of nuanced urban planning policies, fostering locally informed and globally connected approaches to efficiently support urban resilience and sustainability (Elmqvist et al., 2019).
Traditionally, obtaining building attributes has involved expert evaluation, government records, or crowdsourced labeling, which often require field studies. This approach limits coverage and efficiency, leaving many buildings without detailed information. Although platforms like OpenStreetMap (OSM) and government databases now contain diverse urban information, the incompleteness and uneven geographical distribution of global building features hinder their usability across larger regions (Biljecki et al., 2023; Lei et al., 2023; Herfort et al., 2023). With their rapid development, remote sensing-based methods have become a standard approach for extracting building information from aerial and satellite imagery, including attributes such as building height (Wu et al., 2023b; Frantz et al., 2021), and types (Du et al., 2015; Zhao et al., 2019). Remote sensing technologies offer significant advantages, such as large-scale data coverage, reduced dependency on ground surveys, and the ability to monitor urban changes over time with high spatial and temporal resolution. In parallel, machine learning methods that leverage geometric and built environment information have been widely applied to enhance the coverage and accuracy of building data (Roy et al., 2023; Milojevic-Dupont et al., 2023; Lei et al., 2024; Wang et al., 2024c). Despite the advancements, the top-down observation of these technologies poses inherent challenges for object-based building evaluation, as the vertical dimension of buildings holds critical and detailed information that is often difficult to capture.
The emergence of easily accessible Street View Imagery (SVI) has transformed the way buildings are analyzed, providing a ground-level, bottom-up perspective that captures architectural details often obscured in aerial or satellite imagery (Biljecki and Ito, 2021; Gaw et al., 2022; Zhang et al., 2024). Leveraging this capability, numerous studies have integrated deep learning with SVI to extract and profile various building attributes, including height (Yan and Huang, 2022; Fan et al., 2024), type and usage (Kang et al., 2018; Zhao et al., 2021; Ramalingam and Kumar, 2023), architectural style (Lindenthal and Johnson, 2021; Sun et al., 2022b), and facade materials (Xu et al., 2023; Raghu et al., 2023; Chen et al., 2024a). Beyond building profiling, these integrations also support a range of practical applications, including risk assessment (Pelizari et al., 2021; Wang et al., 2021), refinement of 3D building models (Zhang et al., 2021), and building energy efficiency estimation (Sun et al., 2022a; Mayer et al., 2023). These advancements have significantly contributed to SVI-based urban studies, enabling fine-grained, large-scale geospatial analyses.
Despite advancements in SVI-based methods for inferring building attributes, various challenges limit scalability and adaptability: (1) existing datasets struggle with uncertainty due to limited angular coverage in perspective views or distortions in panoramic images, hindering comprehensive observations; (2) reliance on proprietary data restricts accessibility, transparency, and adaptability, with ambiguous licensing further limiting research utility and inclusivity (Helbich et al., 2024); (3) while some efforts align visual data with geolocation, annotations often focus on isolated attributes, requiring separate models. Multi-task learning has been explored (Chen et al., 2022), but class diversity and scalability remain constrained. Rigid annotation schemes further prevent adaptation to emergent characteristics like mixed-use functions or hybrid materials, limiting the ability to capture architectural complexity for more inclusive and interpretable analyses. As a consequence of these challenges, the volume of building datasets derived from SVI that offer a holistic view of structures, fully rely on open datasets, and support comprehensive architectural insights remains quite limited.
Vision-language models (VLMs) — which integrate advanced computer vision (CV) and natural language processing — have demonstrated the ability to interpret complex visual relationships, reason about scenes, and generate coherent, semantically rich descriptions (Li et al., 2024a). In the remote sensing domain, vision-language tasks have demonstrated promise for multi-scale feature understanding, multi-task learning, and applications such as visual question answering, image captioning, and semantic segmentation (Zia et al., 2022; Hu et al., 2023; Dong et al., 2024; Wang et al., 2024a). More recently, multimodal large language models (MLLMs) have advanced these capabilities by integrating deep contextual and semantic representations learned from massive, multimodal datasets, thereby enabling more nuanced and precise interpretations of visual data. This versatility highlights their potential to serve as foundational instruments in SVI-based building research, by enhancing the characterization of building properties, streamlining multi-task learning, and transcending predefined label sets in the analysis of facade features.
To advance fine-grained, bottom-up observations of buildings, we propose an open framework, OpenFACADES, that enriches a variety of building properties from a street-level perspective by leveraging multimodal crowdsourced inputs and open-source MLLMs. First, we utilize fully open-source building footprints and SVI to perform visibility simulations that geospatially align and integrate visible building geometries with corresponding SVI shooting locations. Second, we introduce an innovative pipeline that detects individual buildings based on their visible angles and acquires holistic building images using a custom image reprojection method. Third, we assemble one of the largest global, multi-attribute building image datasets by combining crowdsourced building attributes with high-quality text descriptions generated by state-of-the-art MLLMs. Leveraging this dataset, we are among the first to introduce tailored MLLMs for building profiling through multi-task learning, encompassing both single- and multi-attribute prediction tasks as well as open-vocabulary captioning. Furthermore, we present an in-depth comparative analysis of model performance across various hyperparameter settings, cross-city generalization scenarios, and image quality variations, discussing key challenges identified in prior urban studies (Sun et al., 2022b; Hou et al., 2025; Hou and Biljecki, 2022). These experiments serve as the first systematic demonstration of applying multimodal large language models (MLLMs) in building-level studies, paving the way for their broader adoption in urban research.
The primary contributions of this work are threefold:
  • Developed a reproducible methodology that (1) geolocates, detects, and acquires holistic building images from crowdsourced SVI; (2) integrates these images with crowdsourced building data to create an open and structured building image dataset; and (3) enables future scalability by dynamically retrieving the latest available data from these sources.
  • Compiled an open global building dataset, consisting of (1) 30,180 individual building images from seven cities across three continents, annotated with attribute labels from OSM and text descriptions generated by ChatGPT-4o; and (2) large-scale automated annotations on 1.2 million images covering over half a million buildings. Each image is linked to its geospatial location and enriched with diverse attributes (e.g., building type, number of floors, age, and surface material) along with detailed textual descriptions. This forms the OpenFACADE dataset, one of the largest such resources, spanning multiple urban morphologies.
  • Introduced the first benchmark open-source MLLMs that (1) perform multi-attribute prediction on buildings, achieving robust and more accurate image labeling performance than zero-shot ChatGPT-4o; (2) generate descriptive captions on architectural features, providing comprehensive information beyond standard building attributes; and (3) demonstrate enhanced robustness and generalizability relative to prior CV models.
In summary, this work presents a comprehensive and reproducible framework that leverages multimodal crowdsourced data to develop a global street-level building dataset for training multimodal models. This approach enhances the scope, adaptability, and accuracy of urban analysis, enabling more detailed and interpretable assessments of the built environment.
2 Related work
2.1 Existing street-level building datasets
With advances in geospatial artificial intelligence technologies, research in recent years has increasingly leveraged remote sensing datasets such as high-resolution satellite and aerial imagery, and LiDAR to enhance urban development and planning applications. These datasets enable object-based image analysis, pixel-based classification, and semantic segmentation of urban structures, providing critical insights for land use mapping, urban morphology analysis, and spatiotemporal change detection. As key urban components, buildings have spurred the creation of domain-specific datasets and methodologies to support applications such as urban sustainability evaluation through rooftop attributes extraction (Wu and Biljecki, 2021), infrastructure management via automated land cover classification (Boguszewski et al., 2021), and disaster management through assessing damage (Gupta et al., 2019; Li et al., 2025a).
SVI, rapidly emerging as a prominent proximal remote sensing data source, has been leveraged to generate spatially enriched urban datasets that facilitate fine-grained semantic understanding of complex urban scenes (Biljecki and Ito, 2021). Among these, building-centric SVI datasets enable facade-level feature extraction, offering images that capture textural, material, and architectural features of building exteriors for environmental modeling. Building age and architectural style — key indicators of urban morphological evolution — have long been studied for their correlations with building thermal performance (Tooke et al., 2014; Aksoezen et al., 2015; Nouvel et al., 2017) and influence on pricing models in real estate (Zietz et al., 2008; Lindenthal and Johnson, 2021). Recent advances include the work of Sun et al. (2022b), which applies deep convolutional neural networks (CNNs) to classify buildings in Amsterdam, the Netherlands, into architectural periodization categories (e.g., revival, postwar). Material characterization (Xu et al., 2023; Chen et al., 2024a), another aspect critical for building energy simulation (Nouvel et al., 2017), also supports circular economy objectives by enabling lifecycle material tracking (Raghu et al., 2023) and risk assessment (Wang et al., 2021). Among these efforts, Raghu et al. (2023) employ a multi-city material categories (brick, stucco, etc.) using geotagged SVI perspective views, aligning visual patterns with ground-truth material information for scalable building classification. Combining the aspects of building age and material, Ogawa et al. (2023) introduced a method to detect and geolocate buildings from panoramic images, automatically annotating them with objective building data in Kobe, Japan.
Furthermore, building type or usage — a critical attribute in urban remote sensing and land use classification frameworks — is another important aspect in street-level research (Kang et al., 2018; Zhao et al., 2021; Lindenthal and Johnson, 2021; Ramalingam and Kumar, 2023; Li et al., 2025b). A seminal work by Kang et al. (2018) introduces the BIC_GSV dataset, a multi-city geospatial database of 19,658 SVI-derived building facades categorized into eight classes (e.g., apartment, church, garage, etc.) across North America. These ground truth labels are generated through view-direction-aligned spatial joins with OSM building footprints, enabling parcel-scale urban pattern analysis. Advancing this, Zhao et al. (2021) developed the BEAUTY dataset, which extends BIC_GSV by incorporating both SVI-based land use classification (e.g., residential, commercial, etc.) and multi-class building detection. Other similar research frameworks have also been applied to large-scale urban studies, integrating additional building attributes such as floor number estimation, abandoned house detection, and seismic risk assessment (Iannelli and Dell’Acqua, 2017; Zou and Wang, 2021; Rosenfelder et al., 2021; Pelizari et al., 2021; Ghione et al., 2022). These workflows not only enable location-based building retrieval but also demonstrate cross-modal alignment of SVI with open geospatial building footprints, optimizing multi-source training data generation for CV pipelines.
However, several challenges still remain in street-level building research, limiting the scalability and adaptability of current approaches. First, although many efforts have aligned visual information with building geolocation (Kang et al., 2018; Sun et al., 2022c; Ogawa et al., 2023), they are often either reliant on perspective views with restricted angular coverage, limiting visibility of upper building elements, or on panoramic images prone to severe distortions, misaligning with actual observations. Second, while various SVI-based building datasets have been established, their dependence on data derived from proprietary platforms introduces limitations related to accessibility, transparency, and adaptability. The ambiguous licensing terms of such datasets further constrain their utility for diverse research applications and compromise the integrity of work built upon them, thereby hindering inclusivity within the research community (Helbich et al., 2024). In a recent trend, crowdsourced SVI platforms have garnered attention in urban studies by producing diverse, publicly accessible imagery, prompting efforts to expand dataset coverage and customized applications. Examples include annotating points of interest (Zarbakhsh and McArdle, 2023), image status (Hou et al., 2024), human perception (Yang et al., 2025), and road surface type (Kapp et al., 2025). Among these, Hou et al. (2024) curate a manually labeled dataset to assess 10 million crowdsourced SVIs from 688 cities, enriched with metadata such as platform, weather, and lighting conditions, while Kapp et al. (2025) utilize OSM tags and ChatGPT-4o to label and amplify underrepresented road surface classes, resulting in 9,122 labeled images. These initiatives illustrate the potential of crowdsourced data for broad, inclusive urban analyses.
2.2 Vision models in urban analytics
With the rapid development of deep learning techniques over the past decade, diverse methods have been developed to extract urban cues from visual information, providing efficient and scalable frameworks for understanding built environments. In terms of building facade research, in particular, Convolutional Neural Networks (CNNs) have been widely employed due to their strong feature representation capabilities. Among them, VGG, DenseNet, and ResNet have been extensively applied to achieve, or serve as benchmarks for, the accurate classification and evaluation of building functions (Kang et al., 2018), materials (Ghione et al., 2022; Raghu et al., 2023), architectural styles (Lindenthal and Johnson, 2021; Sun et al., 2022b; Ogawa et al., 2023), and human perceptions (Liang et al., 2024). Additionally, Vision Transformers (ViTs) have emerged as powerful alternatives, leveraging self-attention mechanisms to capture long-range dependencies in building images. Recent studies have demonstrated the effectiveness of ViTs in urban analytics, achieving state-of-the-art performance in material recognition, and construction period prediction (Raghu et al., 2023; Ogawa et al., 2023). Beyond that, hybrid models combining various model backbones have been further developed to consider multi-dimensional features as input, improving comprehensiveness and generalizability in multi-scale urban analysis (Huang et al., 2023; Jia et al., 2024; Fujiwara et al., 2024).
However, the annotation of building attributes remains a fundamental limitation in these approaches. Labels are often restricted to isolated attributes, such as building type or material, necessitating the training of separate models for different objectives. While multi-task learning frameworks have been explored (Chen et al., 2022), class diversity and model scalability remain constrained. Moreover, annotation schemes are typically predefined and rigid due to the availability of data, preventing adaptation to unannotated or emergent building characteristics, such as mixed-use functions or hybrid architectural materials. This lack of multi-dimensional, context-aware labels significantly limits the ability to capture architectural complexity, hindering the development of comprehensive, inclusive, and interpretable approaches for building analysis.
Rapid advancements in LLMs offer new avenues for extracting nuanced insights about complex urban environments. Notably, VLMs combine visual and linguistic modalities, leveraging deep semantic reasoning to establish rich connections between visual concepts and textual descriptions (Wu et al., 2023a; Li et al., 2024a). Building on these capabilities, recent work in remote sensing demonstrates how VLMs can exceed traditional CV methods by producing more context-aware and human-like interpretations (Al Rahhal et al., 2022; Zia et al., 2022; Hu et al., 2023), thereby providing not only precise visual recognition but also a semantic understanding of objects and their relationships within complex environments. In terms of street-level building research, recent studies have explored the state-of-the-art models for automated building annotation. For example, Li et al. (2024b) employed ChatGPT-4o to generate structured multi-label annotations for buildings using SVIs across multiple cities. Similarly, Zeng et al. (2024) assessed the model’s performance in zero-shot building age prediction, finding that ChatGPT-4 effectively estimates the construction period of buildings based on a single-perspective view. However, deploying proprietary LLMs such as ChatGPT-4o at scale presents limitations. Model inference relies on API-based access, which incurs high computational costs, making large-scale applications financially and computationally restrictive, which also constrains the efficiency for fine-tuning, limiting their adaptability for domain-specific urban studies. To address these challenges, recent open-source initiatives have produced diverse series of LLMs, including Qwen-VL (Wang et al., 2024b), Llama (Dubey et al., 2024), and InternVL (Chen et al., 2024b), enabling greater customization and efficiency in downstream tasks. These models exhibit unified capabilities to process multi-dimensional inputs, generating context-aware descriptions informed by their pretraining on large-scale, diverse datasets. By effectively capturing the distributions of natural language and multimodal semantics, open-source VLMs exhibit strong generalization performance while maintaining the flexibility for task-specific adaptations. This capability holds significant potential for advancing street-level urban analysis, as their ability to interpret human-centric observations closely aligns with how individuals perceive and contextualize the built environment.
Hence, we propose a reproducible methodology for integrating open-source multimodal building data from global cities into a comprehensive dataset, incorporating objective attributes and detailed VLM-based interpretations. Table 1 provides an overview of existing SVI datasets related to building attributes, highlighting how our contribution addresses current limitations while significantly expanding the scale, scope, and dimensionality of SVI-based datasets for building-related research. This advancement not only enhances the accessibility and adaptability of building datasets but also paves the way for broader, more inclusive, and scalable applications in urban analytics.
Table 1: Characteristics of existing SVI-based datasets constructed for building-oriented CV and urban research applications, and the features of the dataset we established in this research (GSV: Google Street View).
Studies Purpose Lineage Coverage Category Task Building attribute Image source Image type No. of labeled images No. of cities Continent(s) BIC_GSV (Kang et al., 2018) image classification type GSV perspective 19,658 More than 30 North America apartment, church, garage, house, industrial, office building, retail, roof (8 categories) BEAUTY (Zhao et al., 2021) image classification and object detection type GSV perspective 19,070 More than 30 North America Image classification: residential, commercial, public, industrial (4 categories);
Multi-class detection: apartment, church, garage, house, industrial, office building, retail, roof (8 categories).
Lindenthal and Johnson (2021) image classification age GSV perspective 29,177 1 Europe Georgian, early Victorian, late Victorian/Edwardian, interwar, postwar, contemporary, revival (7 categories). Raghu et al. (2023) image classification surface material GSV perspective 985 3 Asia, North America, Europe brick, stucco, rustication, siding, wood, metal, other (7 categories) SVI4BuildingFunc (Li et al., 2025b) object detection type GSV panoramic 15,400 4 North America, Europe varies by city (e.g., high residential, low residential, commercial, office, walk-up buildings, mixed-up buildings; 5 to 6 categories per city) OpenFACADES Image labeling and captioning type, age, floor, surface material, feature description Mapillary individual building images 30,180 7 North America, Europe, Asia Type: apartments, house, retail, office, hotel, industrial, religious, education, public, garage (10 categories);
Surface material: metal, glass, brick, stone, concrete, wood, plaster (7 categories);
Age: numeric value;
Floor: numeric value.
3 Methodology
In this study, we introduce OpenFACADES, a comprehensive framework for acquiring building images from Street View Imagery (SVI) and automatically annotating them with crowdsourced data, supported by a series of methodological advancements. This framework facilitates the development of large multimodal models tailored for architectural attributes question-answering and captioning. The framework is structured into three main steps, as illustrated in Figure 1:
(1) Integrating multimodal crowdsourced data. Initially, crowdsourced SVI metadata and building data are collected for the designated research areas. Then, isovist analysis is performed to simulate the theoretical angles of view (AOV) from each camera location to the target structures. Based on observation quality, high-quality SVIs are retrieved and further filtered based on their image features, ensuring that only candidate images with optimal visibility and clarity are retained for subsequent analysis.
(2) Retrieving building image data. Based on the geospatial AVOs simulated, we map the relative viewing angles and detect target buildings within the image space. This process enable us to precisely identify associate building information with their visual representations. Then, based on the coordinates of bounding boxes, building images are reprojected from panoramic to perspective view, generating observation aligned with real world observation. These images further undergo a image filtering process to identify high-quality and suitable building images.
(3) Establishing dataset and multimodal models. Building images with available crowdsourced data form a dataset with four label types: categorical, single-word Q&A, multi-attribute Q&A, and captioning. Categorical labels are derived from building information, while single-word Q&A labels append categorical labels to targeted questions, generating concise question-to-label pairs. Multi-attribute Q&A and captioning labels are generated using ChatGPT-4o, enabling detailed textual descriptions and structured annotations for comprehensive building attribute analysis. The last three label types are utilized to fine-tune vision-language models, enabling a versatile model for multi-attribute building labeling and captioning with enhanced contextual understanding.
Figure 1: General framework for integrating multimodal crowdsourced data to establish a street-level building dataset and develop multimodal models. Data: (c) Mapillary and OpenStreetMap contributors.
3.1 Integrating multimodal crowdsourced data
The workflow of integrating multimodal crowdsourced data for building analysis is illustrated in Figure 2. The process includes: (1) preprocessing street-level imagery based on metadata; (2) refining building footprints and attributes; (3) calculating angle of view (AOV) to assess building visibility; and (4) selecting high-quality target images based on AOV thresholds and quality metrics.
Image data preparation
At the first stage, the raw metadata of street-level image data from crowdsourced platform is obtained within study areas before requesting the images. Here, Mapillary is chosen for its extensive global coverage, high-quality user-generated content, and open-access policies that enable reproducible and scalable urban research (Hou and Biljecki, 2022; Kapp et al., 2025; Danish et al., 2025). Specifically, the metadata — comprising location coordinates (computed_geometry), compass angle (computed_compass_angle), capture time (captured_at), and quality indicator (quality_score) — is utilized to structure sorting and quality assessments. Here, filtering operations remove images captured outside the defined study area, exclude multiple images from the same spatial point to prevent redundant viewpoints, and discard those with poor resolution or quality defects. The output of this phase is a curated set of image metadata, with their corresponding unique image IDs, coordinates, and compass angles, prepared for subsequent spatial analyses.
Figure 2: Workflow for obtaining and integrating suitable multimodal crowdsourced data, combining street-level imagery from Mapillary and building information from OpenStreetMap, along with external sources such as Overture Maps and government data, to harmonize building dataset. Data: (c) Mapillary and OpenStreetMap contributors.
Building data preparation
In parallel, building geometries and associated metadata, retrieved from OpenStreetMap (OSM) [^1] https://openstreetmap.org/, are selected as the base dataset for the footprint layer and building information to undergo a refinement process. Based on the context of cities, data harmonization is conducted to supplement missing building footprints from OSM, as well as supplement insufficient building attributes from other data sources, such as Overture Maps [^2] https://overturemaps.org/ and government datasets. Attributes commonly include unique identifiers, building type, facade material, number of floors, construction dates, and polygon geometries. Inconsistencies and outliers — such as footprints representing insignificant or extraneous structures (e.g., roof and underground structures), duplicates introduced by overlapping contributions, or buildings located outside the target region — are systematically removed. After applying these filters, the remaining dataset delivers a precise, consistent, and high-quality representation of the built environment, ready for geometric calculations and alignment with the image data.
Isovist analysis
With both image and building datasets prepared, the next step involves performing isovist analysis to compute theoretical AOV from each camera location to the target structures, building on previous studies (Lindenthal and Johnson, 2021; Ogawa et al., 2023; Fan et al., 2025). This analysis identifies each building’s perimeter segments that fall within the camera’s potential field of view and evaluates the observation efficiency of buildings from specific vantage points. First, a search radius of 50 meters is established to identify surrounding buildings from the SVI capture points. Second, sampling points are generated along the polygonal geometries of buildings within the distance threshold, and lines of sight are constructed towards all sampled points of the target buildings. Third, lines of sight intersecting with surrounding building footprints are filtered out, leaving only the largest angular span between the unobstructed lines, which represents the AOV to a building from a given image shooting point. Additionally, the left and right boundaries of the AOV are recorded as azimuth angles relative to the true north, providing detailed spatial orientation for subsequent tasks. This process identifies which buildings are potentially visible from each image capture point, thereby aligning the building information with the corresponding imagery metadata.
Candidate image selection
Based on the theoretical visibility of buildings, the final stage identifies candidate images most likely to provide reliable and interpretable observations. Criteria derived from the absolute AOV eliminate images taken at excessive distances, those with extreme observation angles (in this study, AOV greater than 120 degrees or smaller than 10 degrees are filtered out), and images that essentially duplicate prior perspectives. Given that crowdsourced SVI, as typical for volunteered geographic information (VGI), can vary in quality and may contain various errors (Hou and Biljecki, 2022), the selected images are then retrieved from Mapillary and evaluated against additional quality metrics — such as brightness, sharpness, and visual complexity — to further identify suitable images for the dataset. Images captured under unsuitable conditions (e.g., nighttime, severe overexposure) or containing excessive visual clutter are removed based on the CV models released in NUS Global Streetscapes (Hou et al., 2024). The result is a high-quality, focused selection of candidate street-level images, optimized for integration with building data in subsequent object detection workflows.
3.2 Retrieving building image data
Figure 3 demonstrates the pipeline for extracting and selecting building images from street-level imagery. The process consists of three main steps: Building detection, image reprojection and image filtering:
Building detection
Azimuth angles derived from isovist analysis are first used to map the relative viewing angles of a building within the image space. This conversion defines a focused AOV for the target building before applying object detection. To determine the position of buildings within panoramic imagery, their relative horizontal coordinate ratios are computed as follows:
$\quad P^{n,i}_{\{l,r\}}=\frac{(A^{n,i}_{\{l,r\}}-H^{i}+C)\mod 360}{360}$ (1)
where $P$, which ranges from 0 to 1, represents the left ($l$) or right ($r$) horizontal coordinates ratio of building $n$ in the panoramic image $i$. The term $H$ denotes the yaw angle when the SVI image token, and $C$ is an adjustable calibration constant that ranges from (0-360), depending on the part of the image the view is oriented towards. Typically, $C$ is set to 180 in Mapillary, indicating that the center of the image is the focal direction.
After determining the relative position of buildings in the SVI, images are cropped using the calculated horizontal coordinate ratios to isolate the AOV focused on the target buildings. Within the focused view, object detection is performed to identify buildings. To accomplish this, we employ GroundingDINO, a model equipped with pre-trained weights capable of detecting various objects using human inputs such as category names or referring expressions (Liu et al., 2023). Specifically, we use the “GroundingDINO-B” model checkpoint, which is trained on several widely-recognized object detection datasets, including COCO, O365, and OpenImage. By assigning the category name “building” to this open-set detector, we generate bounding boxes around the buildings in each cropped image area. This process constrains the observation area to focus on each building footprint, enabling the association of visual observations with 2D building geometries. Additionally, it facilitates the object detection model in isolating target buildings from surrounding elements, such as adjacent structures and environmental noise.
Image reprojection
Panoramic images are formed by mapping the 3D environment onto a 2D sphere, which causes straight lines and familiar shapes to appear curved or distorted. After retrieving the bounding box information from the object detection, the reprojection process is designed to correct these inherent distortions. The objective of the reprojection is to take the portion of the panoramic image identified by the bounding box and present it as if it were photographed by a standard pinhole camera, providing a more intuitive and distortion-free representation of the detected object.
Figure 3: Pipeline demonstrating the extraction and selection of building images from street-level imagery, involving object detection, pixel coordinate transformation and reprojection, and feature-based filtering. Data: (c) Mapillary contributors.
First, we interpret the bounding box region in terms of pixel coordinates within the panoramic imagery, obtaining the box center as ($c_{u}$, $c_{v}$), along with its $width$ and $height$, which are essential for subsequent tasks. Second, a virtual pinhole camera model is constructed based on the specified AOV to a target building and the bounding box $width$. The focal length $f$ and principal point ($c_{x}$, $c_{y}$) in camera coordinate are computed as:
$f=\frac{\frac{width}{2}}{\tan\left(\frac{AOV}{2}\cdot\frac{\pi}{180}\right)}$ (2)
$c_{x}=\frac{width-1}{2},\quad c_{y}=\frac{height-1}{2}$ (3)
These values are used to construct the intrinsic camera matrix $K$, which encapsulates the intrinsic parameters of the virtual pinhole camera. For each pixel ($x$,$y$) in the virtual panel, the transformation from the 2D pixel location to a 3D direction in the camera’s coordinate system is achieved by applying the inverse of $K$:
$K=\begin{bmatrix}f&0&c_{x}\\ 0&f&c_{y}\\ 0&0&1\end{bmatrix},\begin{bmatrix}X{\prime}\\ Y{\prime}\\ Z{\prime}\end{bmatrix}=K^{-1}\begin{bmatrix}x\\ y\\ 1\end{bmatrix}$ (4)
where resulting vector $\mathbf{v}_{cam}=(X{\prime},Y{\prime},Z{\prime})^{T}$ represents the direction of a ray emanating from the camera center through the corresponding pixel on the virtual image plane.
Third, to determine the approximate view direction of the bounding box, we use the center coordinates ($c_{u}$, $c_{v}$) of the bounding box in panoramic coordinate system and combined rotation matrix $R$ to align the camera’s direction to the rotated direction in 3D space:
$\theta=\left(c_{u}-0.5\right)\cdot 360,\quad\phi=\left(0.5-c_{v}\right)\cdot 180$ (5)
$R=R_{x}(\phi)R_{y}(\theta)$ (6)
$\mathbf{v}_{rot}=R\mathbf{v}_{cam}$ (7)
where $c_{u}$ and $c_{v}$ are normalized to a range of [0, 1], with $c_{u}$ as the horizontal center and $c_{v}$ as the vertical center of the bounding box region. The yaw angle $\theta$ defines the horizontal rotation of the camera and spans from $-180^{\circ}$ to $180^{\circ}$. The pitch angle $\phi$ defines the vertical rotation of the camera and spans from $-90^{\circ}$ to $90^{\circ}$. The combined rotation matrix $R$ is formed as the product of two individual rotation matrices based on Rodrigues’ formula: $R_{y}$($\theta$), which rotates the coordinate system around the y-axis (yaw), and $R_{x}$($\phi$), which rotates the coordinate system around the x-axis (pitch). $\mathbf{v}_{cam}$ is the original direction vector in the camera’s coordinate system, while $\mathbf{v}_{rot}$ is the new direction vector after applying the rotations, pointing toward the desired region of the spherical panorama.
Lastly, the rotated 3D direction vector $\mathbf{v}_{rot}=(X,Y,Z)$ is normalized and converted into spherical coordinates, where longitude $\lambda$ and latitude $\varphi$ are calculated based on:
$\lambda=\arctan 2(X,Z),\quad\varphi=\arcsin(Y)$ (8)
The corresponding pixel coordinates ($X_{\text{img}}$, $Y_{\text{img}}$) in the original panoramic image (equirectangular format) are then derived as:
$X_{img}=\left(\frac{\lambda}{2\pi}+0.5\right)(W_{\text{pano}}-1),\quad Y_{img}=\left(\frac{\varphi}{\pi}+0.5\right)(H_{\text{pano}}-1).$ (9)
At these coordinates, pixel values are sampled from the original panoramic image, and reprojected to generate the rectified perspective view using the remap function from OpenCV library. This transformation eliminates the spherical distortions inherent in panoramic imagery, producing a visually intuitive and geometrically corrected view aligned with the detected object. As examples demonstrated in Figure 4, this correction is crucial not only for preserving essential structural details for model interpretation but also for mitigating distortions that could otherwise misalign architectural features. This preprocessing step enhances the model’s ability to accurately analyze building attributes in urban applications.
Figure 4: Examples of different types of building images used as input to the vision-language model, resulting in varied responses. By generating a holistic view of individual buildings, our method facilitates a more authentic analysis and interpretation. Data: (c) Mapillary contributors.
Image filtering
The features of the detected individual building images are further analyzed to refine and enhance the image dataset. ZenSVI (Ito et al., 2024), an open-source library for street-level imagery analysis, is integrated into the framework to facilitate the extraction of image features. We analyze image features across four key dimensions: blurriness, brightness, semantic segmentation, and scene classification. These dimensions are utilized to identify high-quality and suitable building images for inclusion in the dataset: blurriness, brightness, semantic segmentation, and scene classification. Blurriness is evaluated using the OpenCV Laplacian operator to filter out images with motion blur or poor focus, while brightness assessment removes those with suboptimal illumination. A pre-trained Place365 model (Zhou et al., 2017) excludes indoor scenes, and semantic segmentation is applied to detect and minimize occlusions (e.g., trees, vehicles, walls), ensuring that selected images predominantly showcase building facades and maintain high visual quality.
3.3 Establishing dataset and multimodal models
Street-level building dataset
Following the previous process, building information is assigned to the detected buildings in the imagery. In this study, we focus on the classification of building type, facade material, construction year (age), and number of floors, which have been identified as essential attributes in prior studies and are supported by relatively sufficient data for model training and evaluation. Specifically, we utilize labels from building data corresponding to the categories: building, building:material, start_date, and building:levels, which are primarily sourced from OSM and supplemented by additional building datasets, as mentioned in Section 3.1. Here, building type and facade material are treated as categorical variables, while construction year and number of floors are represented as numerical values.
From the full set of building data, we sample buildings with available category labels to construct the dataset for subsequent model development. The dataset is assembled and divided into training and test sets based on the following three principles: (1) ensuring sufficient labels across all classes to avoid biased predictive accuracy in machine learning tasks, and balanced class distribution in test set; (2) maintaining a balanced geospatial distribution across cities to represent the diversity of architectural designs; and (3) preventing the same building from appearing in both the training and test sets to minimize data leakage.
The dataset contains four types of labels: categorical, single-word Q&A, multi-attribute Q&A, and captioning labels. Categorical labels are used to fine-tune CNN/ViT-based vision models, serving as the baseline for evaluating the performance of common practices. Single-word Q&A labels are derived from categorical labels by appending the label to specific questions about the four building attributes, thereby generating concise question-to-label pairs based on building information. Multi-attribute Q&A and captioning labels are generated using the state-of-the-art multimodal large language model, ChatGPT-4o, through the OpenAI API [^3] https://www.openai.com/. This task involves prompting the model to annotate or describe the building features visible in the images, thereby creating an image-text training set. Table 2 provides detailed indication of data sources and examples of these labels, showcasing the comprehensive structure of the dataset designed to facilitate the interpretation of building imagery.
Table 2: Different label types and data collection approaches for developing a street-level building dataset.
Label type Target model Data source Examples
Categorical label Computer vision models OpenStreetMap + government data apartment, house, retail, etc.
Single-word Q&A Vision-Language models OpenStreetMap + government data Q: Assign a brief label for the floor count of the building in the image. A: 12
Multi-attribute Q&A Teacher model (ChatGPT-4o) Q: Provide concise labels for each category using the following JSON format. Select appropriate values from the provided options for each category: {“building_type”: “(choose one option from: ‘apartments’, ‘house’, ‘retail’, ‘office’, ‘hotel’, ‘industrial’, ‘religious’, ‘education’, ‘public’, ‘garage’)”, “alternate_building_type”: “(choose one option from: ‘apartments’, ‘house’, ‘retail’, ‘office’, ‘hotel’, ‘industrial’, ‘religious’, ‘education’, ‘public’, ‘garage’)”, “building_age”: “(a 4-digit year indicating the approximate construction date of the building)”, “floors”: “(a numeric value representing the total number of floors)”, “surface_material”: “(choose one option from: ‘brick’, ‘wood’, ‘concrete’, ‘metal’, ‘stone’, ‘glass’, ‘plaster’)”, “alternate_surface_material”: “(choose one option from: ‘brick’, ‘wood’, ‘concrete’, ‘metal’, ‘stone’, ‘glass’, ‘plaster’)”} A: {“building_type”: “office”, “alternate_building_type”: “retail”, “building_age”: 1950, “floors”: 8, “surface_material”: “brick”, “alternate_surface_material”: “stone”}
Captioning Teacher model (ChatGPT-4o) Q: Analyze the building shown in the image and provide a detailed description of its architectural features. Then, describe the building type, the building’s age (by specifying an approximate construction year), the primary facade material (the main material visible on the building’s surface), the construction material, and the total number of floors in the building. A: The building in the image appears to be a residential structure, likely a house …… In summary, the building is a traditional wooden house with a steeply pitched roof, likely constructed in the late 19th to early 20th century. It has a single-story design with a wooden facade and is well-maintained.
Vision-language models
To address the limitations of traditional categorical classification models in building attribute analysis, we leverage InternVL2.5 (Chen et al., 2024b), an open-source multimodal large language model (MLLM) designed for unified visual-language reasoning. As depicted in Figure 5, InternVL2.5 is built on the “ViT-MLP-LLM” paradigm by integrating a scalable vision encoder (InternViT) (Chen et al., 2024c), a multi-layer perceptron (MLP) projector, and a large language model (LLM). The vision encoder is InternViT-300M-448px-V2.5, a distilled variant of the 6B-parameter model optimized via dynamic high-resolution training and next token prediction (NTP) loss (Chen et al., 2024b). This architecture processes 448×448 pixel image tiles through a pixel unshuffle operation, reducing 1024 visual tokens to 256 for efficient cross-modal alignment.
The model is selected for its general-purpose captioning and open-vocabulary classification capabilities, critical for capturing the multifaceted attributes of buildings (e.g., material, style, type) within a unified framework. Unlike conventional models restricted to predefined labels, InternVL2.5’s contrastive vision-language pretraining enables semantic reasoning over diverse facade characteristics, aligning with our goal of holistic building profiling. Full-model tuning is conducted through optimizing three components (Figure 5): (1) InternViT-300M Vision Encoder: Retrained on street-level building images to enhance facade feature extraction, leveraging dynamic high-resolution (448px) inputs; (2) MLP Projector: Adjusted to align building-specific visual tokens with textual embeddings in the LLM space; (3) LLM Head: Fine-tuned using the corpus of building characteristic descriptions to generate structured captions.
Figure 5: The overall framework of the InternVL series model architecture for building-centric tasks. Data: (c) Mapillary contributors.
After fine-tuning, we design a comprehensive set of experiments to evaluate the model’s performance in several respects: its general capability of VLMs, its sensitivity to training data size (ablation analysis), and its effectiveness in inferring building attributes across different cities and categories. To align these investigations more closely with real-world building profiling practices, we further conduct comparative assessments involving both VLMs and conventional categorical classification models. These comparisons target three main objectives: (1) examining the performance of the fine-tuned models, (2) assessing their generalizability to unseen data (a crucial concern in cross-city studies, where models trained on one urban area may not perform consistently when applied elsewhere (Sun et al., 2022b; Hou et al., 2025)), and (3) evaluating robustness to heterogeneous noises and degradation. The latter is particularly relevant because crowdsourced images, unlike standardized remote sensing imagery, frequently exhibit diverse quality issues (Hou and Biljecki, 2022). These experiments offer deeper insights into models’ generalization capacity.
To investigate robustness and impact of common issues in imagery, we adopt the methodology outlined by Hendrycks and Dietterich (2019), which measures model resilience against common image corruptions and perturbations. Guided by the image quality criteria proposed by Hou and Biljecki (2022), we algorithmically generate four types of corruptions (Figure 6) — occlusion, motion blur, Gaussian noise, and brightness alterations — and apply them to the test set. We then evaluate each model’s performance under these degraded conditions using Relative Corruption Errors ($Relative$ CE) (Hendrycks and Dietterich, 2019). First, the baseline error rate $E^{m}_{clean}$ was determined for model $m$ on the uncorrupted data. Next, we compute the error rate $E^{m}_{c,s}$ for each corruption type $c$ at severity level $s$ (1 $\leq$ s $\leq$ 3). In classification tasks (building type and surface material), the error rate is defined as $1-Accuracy$, whereas for regression tasks (predicting number of floors and building age), it is defined as $1-R^{2}$. Finally, to account for the varying difficulties introduced by each corruption, we normalize these error rates by dividing by the ResNet50 baseline error; $Relative$ CE is calculated as:
$RelativeCE^{f}_{c}=(\sum^{3}_{s=1}E^{f}_{s,c}-E^{f}_{clean})/(\sum^{3}_{s=1}E^{ResNet50}_{s,c}-E^{ResNet50}_{clean})$ (10)
This normalization provides a clearer measure of how much each model’s performance declines under different corruptions. Averaging these $Relative$ CE from four types of corruptions results in the $Relative$ $m$CE, which represents the overall relative performance degradation when the models encountering corruptions.
Figure 6: Examples of image corruption and perturbation for robustness experiments, consisting of four categories of algorithmically generated images based on common quality issues in crowdsourced imagery. Each type of corruption has 3 levels of severity (except for brightness which has twice 3 levels of severity), resulting in a total of 15 corruption levels. Data: (c) Mapillary contributors.
4 Experiments and results
4.1 Street-level building dataset
The building dataset is established using panoramic images sourced from Mapillary [^4] https://www.mapillary.com. We manually select cities that have a sufficient number of panoramic images available through the Mapillary online interface, and that also have a considerable amount of objective building attributes openly in OSM. Ultimately, seven cities from three continents are chosen, including Amsterdam, Berlin, Helsinki, San Francisco, Washington D.C., Houston and Manila, balancing the dataset across both selection aspects. Among them, Helsinki is selected due to its rich availability of building material data from the Buildings in Helsinki data [^5] https://hri.fi/data/en_GB/dataset/helsingin-rakennukset, while Amsterdam provides diverse data on building age, to add sufficient data on according aspects.
The metadata for panoramic SVIs is first downloaded within the defined city boundaries using the Mapillary Python Software Development Kit [^6] https://github.com/mapillary/mapillary-python-sdk, while building data is retrieved using OSMnx (Boeing, 2017). Subsequently, the data undergoes the process described in Section 3.1 to calculate the angle of view, evaluate observation quality, and identify candidate images. These selected images are then utilized for building detection, image reprojection, and filtering, as detailed in Section 3.2, resulting in a collection of individual building images for each city. Table 3 provides a detailed breakdown of the number of buildings, SVI images retrieved, individual buildings detected with associated images, and the ratio of completeness for each city. While completeness varies among cities due to differences in the availability and quality of Mapillary images uploaded for specific locations, around 50% of buildings in city centers can be observed and analyzed.
Table 3: Summary of building footprints, image retrieval, and detection completeness across cities and regions in the building dataset.
City Total building footprints Total images retrieved Total individual building images Buildings with images Percentage detected City center completeness (2.5km×$\times$2.5km)
Europe
Amsterdam 195,188 203,570 330,235 120,154 61.6% 83.6%
Helsinki 63,972 20,035 20,479 8,930 14.0% 42.5%
Berlin 497,703 408,166 287,065 137,930 27.7% 46.7%
North America
San Francisco 160,659 62,521 91,874 34,510 21.5% 39.4%
Houston 399,883 304,030 238,934 91,774 23.0% 53.8%
Washington D.C. 161,190 269,420 201,955 86,144 53.4% 57.4%
Asia
Manila 105,904 68,706 48,951 23,911 22.6% 22.7%
Total 1,617,019 1,414,288 1,219,493 503,353 31.1% 49.4%
As discussed in Section 3.3, building images with available attributes are sampled to construct a class-sufficient dataset for model development, resulting in a total of 30,180 images. Figure 7 illustrates the distribution of images across relevant categories for each attribute, comprising 17,530 images for building type, 2,871 for surface material, 7,228 for floors, and 5,927 for age. The dataset is divided into training, validation, and test sets in a 6:1:3 ratio. For the training and test sets, ChatGPT-4o is employed to generate multi-attribute Q&A and captioning labels, enriching the dataset with additional descriptive annotations for both training and benchmarking. To comprehensively capture ambiguous architectural features, we also obtain an alternative classifier by using the prompts “alternate_building_type” and “alternate_surface_material”, which represent the top-two predictions of the MLLMs.
We acknowledge that the current dataset has limitations, particularly in terms of geographic diversity across continents and the availability of data for certain attributes, such as surface material and building age. Nevertheless, to the best of our knowledge, this dataset is both large and comprehensive compared to previous efforts highlighted in Section 2. Additionally, the reproducible framework established in this study enables future expansion of the dataset as more building images and their associated attributes become available through crowdsourced platforms. This iterative refinement could progressively enhance the dataset’s scope and utility for broader applications.
Figure 7: The distribution of the building images categorized by objective building attributes—type, age, floor, and surface material—selected for each city in dataset.
4.2 Vision-language model evaluation
4.2.1 General performance
As validation, we benchmark our fine-tuned VLMs against various baselines, including ChatGPT-4o and the InternVL2.5 family of models (zero-shot), to evaluate their effectiveness in building-related tasks. The same hyperparameters are applied to fine-tune each VLM, with the number of epochs set to 3 and a learning rate of 8e-6 in a full fine-tuning setup. The results, consolidated in Tables 4(a) and 4(b), confirm the advantages of employing VLMs in these diverse tasks by matching or outperforming ChatGPT-4o baselines in most scenarios.
Table 4: Validation performance comparison of large language models on prediction tasks: (a) building type and surface material; (b) building floors and building age.
(a) Building type and surface material classification in zero-shot and fine-tuned settings.
Attribute Model Size Accuracy Precision Recall F1 Acc@2
Building type Zero-shot
ChatGPT-4o - 0.58 0.64 0.58 0.57 0.75
InternVL2.5 1B 0.44 0.60 0.44 0.42 0.54
2B 0.46 0.56 0.46 0.44 0.51
4B 0.48 0.59 0.48 0.47 0.63
Fine-tuned
InternVL2.5 1B 0.60 0.65 0.60 0.59 0.75
2B 0.61 0.64 0.61 0.60 0.76
4B 0.62 0.66 0.62 0.62 0.77
Surface material Zero-shot
ChatGPT-4o - 0.65 0.70 0.65 0.64 0.79
InternVL2.5 1B 0.59 0.62 0.59 0.58 0.69
2B 0.60 0.63 0.60 0.60 0.72
4B 0.61 0.65 0.61 0.61 0.76
Fine-tuned
InternVL2.5 1B 0.69 0.75 0.69 0.69 0.82
2B 0.69 0.74 0.69 0.68 0.82
4B 0.69 0.74 0.69 0.68 0.82
(b) Building floors and building age prediction in zero-shot and fine-tuned settings.
Attribute Model Size R2 ($\uparrow$) MAE ($\downarrow$) MAPE ($\downarrow$) RMSE ($\downarrow$)
Building floors Zero-shot
ChatGPT-4o - 0.72 2.36 0.39 5.01
InternVL2.5 1B -0.02 5.46 0.59 9.58
2B 0.24 4.74 0.49 8.26
4B 0.55 3.68 0.44 6.53
Fine-tuned
InternVL2.5 1B 0.75 2.26 0.35 4.72
2B 0.77 2.32 0.36 4.53
4B 0.78 2.22 0.35 4.45
Building age Zero-shot
ChatGPT-4o - 0.65 31.63 0.74 57.07
InternVL2.5 1B 0.35 53.09 0.99 78.94
2B 0.31 52.94 2.02 79.35
4B 0.24 51.93 1.05 84.12
Fine-tuned
InternVL2.5 1B 0.70 29.22 0.64 52.35
2B 0.70 29.24 0.66 52.03
4B 0.71 29.11 0.64 51.85
In classification tasks, building type and surface material performances are assessed using accuracy, F1-score, recall, and precision. Fine-tuned VLMs exhibit competitive performance, achieving a Top-2 classification accuracy of 75% for building type and 82% for surface material. For predicting the number of building floors and building age, we employ R-squared, Mean Absolute Error (MAE), Mean Absolute Percentage Error (MAPE), and Root Mean Squared Error (RMSE) as evaluation metrics. In building floor prediction, the fine-tuned InternVL2.5 series of models achieve a high R-squared value of approximately 0.77, an MAE of 2.3 floors, and an RMSE of around 4.5 floors, demonstrating substantial improvements over zero-shot baselines. Similarly, in building age prediction, InternVL2.5 models deliver strong results, with an R-squared of approximately 0.70 and an MAE of 29 years, highlighting their effectiveness in capturing architectural and temporal characteristics.
Table 5: METEOR and ROUGE-L evaluation in zero-shot and fine-tuned settings.
Model Size METEOR ROUGE-L
Zero-shot
InternVL2.5 1B 33.14 26.92
2B 33.06 28.90
4B 34.24 29.38
Fine-tuned
InternVL2.5 1B 39.05 39.06
2B 40.10 39.76
4B 39.17 39.27
Furthermore, the text generated by ChatGPT-4o is used as a baseline to evaluate performance gains in image captioning. In this study, we adopt the METEOR and ROUGE_L metrics, as they assess semantic similarity and linguistic coherence, making them more suitable for evaluating architectural characteristics. Table 5 presents the evaluation metrics across different models. Similar to the labeling task, pre-trained models demonstrate captioning capabilities compared to ChatGPT-4o, while fine-tuned models acquire additional domain-specific knowledge. We observe that increasing model size generally improves performance; however, once the parameter count doubles, marginal gains diminish. This effect is likely due to inherent noise in OSM-derived ground-truth labels and the limited scale of the dataset, which may reduce the advantages of larger architectures. Balancing computational cost with predictive accuracy, we select the 2B-parameter model for subsequent experiments, as it allows efficient inference on a single 24 GB GPU while still delivering notable performance improvements.
4.2.2 Performance by cities
Figure 8 presents the performance of three model variants — (1) the InternVL2.5‐2B model before fine‐tuning and (2) after fine‐tuning, as well as (3) a ChatGPT‐4o reference baseline — on seven cities and four building attributes. In general, the fine‐tuned InternVL2.5‐2B outperforms its non‐fine‐tuned counterpart, showing consistent gains in classification accuracy (Acc.) for building type and surface material, as well as higher R-squared for predicting building floors and age. These improvements are particularly notable in Berlin and San Francisco, where building material, floor and age performance improve substantially. Amsterdam and Helsinki also exhibit moderate but still positive gains for different tasks.
Despite the overall upward trend, improvement magnitude varies across cities and attributes, which may due to several reasons. First, the availability of diverse and distinct samples plays a crucial role: cities with a richer variety of building facades (e.g., Amsterdam, Berlin) yield more pronounced performance boosts. Conversely, locations with more homogeneous or ambiguous building styles (e.g., Manila, Helsinki) show relatively smaller gains. Second, crowdsourced labels in certain cities may be incorrect or insufficient, which can adversely affect the model’s ability to learn reliable city‐specific patterns, restraining potential performance gains.
Nevertheless, when benchmarked against the ChatGPT‐4o baseline, a robust reference point due to its extensive multimodal pretraining, the fine‐tuned InternVL2.5‐2B model demonstrates generally competitive or superior performance. These results confirm that open‐access vision–language models, exemplified by InternVL2.5‐2B, can achieve near‐state‐of‐the‐art performance at no additional licensing cost once adequately fine‐tuned on relevant datasets. This highlights the effectiveness of VLMs in traditional prediction tasks for multiple building attributes across global cities, providing a cost‐effective solution for a wide range of urban remote sensing applications.
Figure 8: Model performance on building attributes across different cities before and after fine-tuning the InternVL2.5-2B model, compared to the baseline performance of ChatGPT-4o. Building type and surface material are evaluated using classification accuracy (Acc.), while number of floors and building age are assessed using R-squared (R2). “NA” indicates cities with insufficient data for model evaluation (ground‐truth instances fewer than 20 in test set).
4.2.3 Performance by categories
Figure 9 presents the confusion matrices illustrating the performance of our VLM on building type and surface material on different categories. Overall, the model demonstrates robust performance for most classes. In terms of well‐predicted labels, visually distinctive building types such as apartments and houses show consistently high accuracies. These categories often have defining features (e.g., apartment blocks characterized by uniform facades and repetitive windows) that the model effectively captures. Similarly, for surface material, high‐frequency and visually salient classes like brick, wood, and glass yield strong performances. Conversely, certain labels are harder to classify, yielding relatively lower accuracies. For building type, hotel or public categories are frequently misidentified as office, suggesting significant overlap in their architectural appearance (e.g., multi-stories, institutional buildings). Likewise, plaster and concrete exhibit misclassifications due to shared grayscale tones and blank textures.
Figure 9: Confusion matrices illustrating the performance of the InternVL2.5-2B model on classifying different categories of building type and surface material. Darker cells indicate higher prediction accuracy.
Figure 10 illustrates the model’s ability to predict building floors and building age under two evaluation schemes: detailed class matching (left) and general range matching (right). For number of floors, the confusion values indicate a good performance for various classes, while when the number increases beyond four or five, misclassification becomes more pronounced. In particular, tall buildings tend to overlap with adjacent categories, highlighting the difficulty of accurately distinguishing high-rise structures based solely on external features and single observation points. When evaluated under the general range scheme, performance improves significantly, suggesting that the model is capable of capturing overall floor patterns from images.
Figure 10: Confusion matrices illustrating the performance of the InternVL2.5-2B model on predicting the number of floors and building age, evaluated based on the accuracy of predictions falling within specific (left) and general (right) ranges. Darker cells indicate higher prediction accuracy.
For building age, the left matrix (exact matching) indicates that categories representing buildings constructed after 1900 exhibit comparatively better performance. This is likely due to their more distinctive architectural styles, which the model can more easily differentiate. In contrast, older buildings (e.g., before 1700, 1800–1900) show greater confusion both among themselves and with intermediate categories (e.g., 1900–1950). This may be attributed to subtle external differences among buildings from these periods and the impact of renovations, additions, or retro designs, which can obscure their original architectural cues (Sun et al., 2022b). Furthermore, the model’s performance may be also influenced by the limited representation of these older eras in the original VLM training dataset.
Taken together, these confusion matrices suggest that the model is capable of inferring the number of floors and the era of construction. However, inherent visual ambiguities — particularly among structurally or stylistically similar categories — contribute to overlaps in predictions. These confusions often arise from overlapping visual cues, which may be further exacerbated by dataset biases, such as incorrect labeling, insufficient sample representation for certain categories, or poor observation. Enhancing the quality, diversity, and coverage of crowdsourced data would be a valuable step toward improving the dataset and the model’s performance.
4.2.4 Ablation experiments on data size
Figure 11 presents the results of ablation experiments on the InternVL2.5-2B using different combinations of a training set from 58,942 image-text pairs. The left panel shows accuracy for material and type predictions, while the right panel illustrates R-squared for floor and age predictions across varying dataset percentages. Table 6 summarizes semantic similarity and linguistic coherence relative to GPT‐generated captions under different conditions.
Figure 11: Model performance across varying dataset sizes by adjusting full training data (solid line) and GPT-generated data (dashed lines). The left plot shows accuracy for building type and surface material classification, while the right plot presents R-squared values for floor and age predictions, benchmarked against ChatGPT baselines.
These ablation experiments reveal several important insights. First, in multi-attribute prediction tasks (Figure 11), performance peaks early in both scenarios of adjusting either OSM data or GPT-generated data. Even smaller datasets (around 5–10% of the full corpus) yield notable performance gains, highlighting the VLM’s ability to learn effectively in data-constrained scenarios. This behavior can be attributed to the pre-trained semantic relationships embedded in the VLM’s latent space from its foundational training. Fine-tuning on limited data stabilizes outputs by aligning task-specific features with the model’s pre-existing knowledge distribution. Second, performance rises gradually when adding OSM ground truth data for most attribute prediction tasks, while GPT-generated data slightly diminishes performance gains. One plausible explanation is that OSM data encodes structured, human‐validated geographic knowledge, whereas GPT‐generated samples may introduce inaccuracies or hallucinated features that misalign the VLM’s learned representations. Mitigating such noise — by refining annotation procedures or excluding low‐quality samples — could improve overall accuracy and robustness (Chen et al., 2024b). Third, adding OSM data constrains the descriptive ability across tasks (Table 6). This trade-off may reflect task interference in multi-task learning: optimizing for structured building attributes could suppress the model’s capacity to generate diverse captions. Addressing this may involve upscaling model capacity, curating high‐quality OSM–GPT hybrid datasets, or leveraging techniques such as knowledge distillation to balance structure with generative expressiveness.
Table 6: METEOR and ROUGE-L evaluation across OSM and GPT splits.
Dataset METEOR ROUGE-L
OSM GPT
- - 33.09 28.97
5% 100% 42.46 41.82
10% 100% 42.02 41.42
25% 100% 41.51 41.00
50% 100% 40.47 40.20
75% 100% 40.81 40.32
100% 5% 34.98 35.72
100% 10% 35.86 36.18
100% 25% 38.25 38.20
100% 50% 39.14 38.83
100% 75% 39.58 39.33
100% 100% 40.28 39.87
In summary, for all structured variables, these results underscore a practical trade‐off: using only 5–10% of the OSM data (1,000-2,000 image–text pairs in this case) can achieve performance superior to the GPT‐4o baseline, while supplementing sufficient OSM data with minimal GPT-generated content can maximize model performance, thereby offering significant efficiency gains. Nonetheless, attaining state‐of‐the‐art descriptive captions likely requires more extensive textual annotations and larger model architectures to capture finer‐grained semantic and linguistic details. Striking a balance between captioning and labeling performance, we ultimately employ the model trained on the full dataset.
4.3 Comparative experiments
4.3.1 Computer vision models
As discussed in Section 2.2, CNNs and ViTs are widely used to infer building attributes. Here, we compare their performance with that of the fine‐tuned InternVL2.5‐2B VLM on four building characteristics, as summarized in Tables 7(a) and 7(b). To facilitate a fair comparison, we integrate GPT-generated data to supplement the missing OSM data in the training set for the CV models.
Table 7: Validation performance comparison between fine-tuned InternVL2.5-2B and CV models.
(a) Performance on classification tasks of building type and surface material
Attribute Model Accuracy Precision Recall F1 Acc@2
Building type DenseNet 0.54 0.54 0.54 0.53 0.68
VGG 0.47 0.47 0.47 0.47 0.64
ResNet 0.52 0.54 0.52 0.52 0.67
ResNet18 0.51 0.51 0.51 0.51 0.67
ResNet101 0.54 0.55 0.54 0.53 0.67
ViT16 0.54 0.56 0.54 0.54 0.69
ViT32 0.52 0.52 0.52 0.51 0.68
InternVL2.5-2B 0.61 0.64 0.61 0.60 0.76
Surface material DenseNet 0.65 0.67 0.65 0.64 0.81
VGG 0.57 0.61 0.57 0.57 0.74
ResNet 0.65 0.67 0.65 0.64 0.79
ResNet18 0.61 0.64 0.61 0.61 0.79
ResNet101 0.66 0.69 0.66 0.65 0.81
ViT16 0.65 0.68 0.65 0.65 0.79
ViT32 0.63 0.67 0.63 0.64 0.78
InternVL2.5-2B 0.69 0.74 0.69 0.68 0.82
(b) Performance on prediction tasks of number of floors and building age.
Attribute Model R2 ($\uparrow$) MAE ($\downarrow$) MAPE ($\downarrow$) RMSE ($\downarrow$)
Number of floors DenseNet 0.77 2.35 0.40 4.27
VGG 0.67 3.15 0.42 5.46
ResNet50 0.78 2.44 0.40 4.54
ResNet18 0.75 2.77 0.40 4.72
ResNet101 0.77 2.52 0.41 4.60
ViT16 0.77 2.44 0.38 4.52
ViT32 0.75 2.58 0.45 4.72
InternVL2.5-2B 0.77 2.22 0.35 4.53
Building age DenseNet 0.72 30.80 1.03 50.26
VGG 0.56 41.65 1.35 63.13
ResNet50 0.72 32.34 1.10 50.14
ResNet18 0.68 34.66 1.11 53.50
ResNet101 0.72 31.25 1.01 50.04
ViT16 0.71 31.82 1.20 50.82
ViT32 0.68 34.45 1.25 53.51
InternVL2.5-2B 0.70 28.64 0.65 51.67
In general, the fine-tuned VLM (InternVL2.5‐2B) achieves the best performance, particularly in tasks such as building type and surface material predictions. While CV models slightly exceed the performance of VLMs in predicting the number of floors and building age when trained on ChatGPT-generated data, these models require domain-specific tuning and separate architectures for each attribute. VLMs, however, offer a unified and highly adaptable approach, achieving comparable or superior performance across multi-attribute prediction tasks. This underscores the advantage of leveraging pretrained semantic reasoning and contextual understanding inherent in VLMs, which generalize well across diverse tasks.
Additionally, VLMs provide the capability to generate textual descriptions of architectural and contextual details, offering richer insights into building attributes and enabling further qualitative or descriptive analyses. This combination of robust predictive performance and expanded functionality makes VLMs a compelling alternative to traditional approaches, particularly in applications requiring adaptability and scalability across complex urban environments.
4.3.2 Generalizability
Generalizing CV models to unseen cities remains a significant challenge due to the diverse and unique architectural features across cities (Sun et al., 2022b). VLMs, pretrained on extensive, high‐quality image‐text datasets, demonstrate promising potential to overcome these limitations by leveraging their pre‐acquired semantic reasoning and contextual understanding capabilities. To investigate this potential, we conducted an experiment on building imagery collected from Brussels, comparing the performance of established CNN and ViT architectures against our fine‐tuned VLM. For this evaluation, we curated a dataset of 3,728 labeled building images by integrating OSM attributes with buildings detected from Mapillary SVI. The dataset is composed of 3,383 images for building type, 186 for surface material, 1,245 for the number of floors, and 99 for building age.
Table 8(a) and 8(b) indicate that the VLM model demonstrates superior generalizability compared to commonly used CV models, which is particularly evident in the tasks of building type and age prediction. The enhanced performance can be attributed to the pre-trained VLM’s ability to leverage semantic reasoning and contextual understanding from its large-scale image-text pretrain data, enabling it to adapt to diverse architectural features. While performance is similar for number of floors predictions, the lower accuracy of ResNet101 and ViT16 on other attributes highlights their reliance on localized visual features, limiting generalization. For surface material classification, VLM achieves high precision, but overall performance is affected by label inconsistencies in the dataset. For instance, buildings labeled as “brick” in OSM are often visually identified as “plaster” during manual inspection, a discrepancy linked to local labeling conventions in Brussels. This issue reflects broader challenges with data quality and the discrepancies in labeling schemes across cities. Developing an unified or context-aware annotation system would be valuable for addressing such inconsistencies.
Table 8: Validation performance comparison of different models on building images in Brussels.
(a) Performance on classification tasks of building type and surface material
Attribute Model Accuracy F1 Precision Recall
Building type ResNet101 0.37 0.46 0.64 0.37
ViT16 0.26 0.35 0.68 0.26
InternVL2.5-2B 0.58 0.63 0.74 0.58
Surface material ResNet101 0.19 0.31 0.93 0.19
ViT16 0.20 0.32 0.84 0.20
InternVL2.5-2B 0.31 0.44 0.93 0.31
(b) Performance on prediction tasks of number of floors and building age
Attribute Model RMSE ($\uparrow$) MAE ($\downarrow$) MAPE ($\downarrow$) R2 ($\uparrow$)
Number of floors ResNet101 1.91 1.12 0.32 0.57
ViT16 1.78 1.06 0.29 0.63
InternVL2.5-2B 1.88 0.94 0.27 0.59
Building age ResNet101 50.72 39.05 1.81 0.15
ViT16 49.36 36.81 1.93 0.22
InternVL2.5-2B 46.03 31.66 1.50 0.44
4.3.3 Sensitivity and robustness
As described in Section 3.3, we evaluate the robustness of the VLM against image corruptions by testing it on a perturbed dataset derived from the test set. Figure 12 illustrates the model’s performance under varying severity levels of occlusion, motion blur, Gaussian noise, and brightness distortions. In general, the model demonstrates resilience, with performance dropping by less than 10% under most mild and moderate image corruptions. In particular, the model remains significantly stable in handling lighting variations and occlusion, both of which are common challenges in crowdsourced image datasets. However, the model experiences a significant performance drop when confronted with moderate to severe noise and blurriness. In particular, the model is most affected by motion blur, where the error rate for the number of floors prediction increases from 0.25 (clean error) to 0.93, and the error rate for building age prediction rises from 0.29 (clean error) to 1.48. These findings emphasize the necessity of image preprocessing techniques to filter out degraded images during the image selection stage.
Figure 12: Error rates of the VLM under different severity levels of image corruption. The dotted line represents the clean error obtained from the original test set, serving as a baseline for comparison.
Furthermore, the Relative $m$CE is computed for all models using ResNet50 as the baseline. Table 9 presents the relative error rates across different building attribute prediction tasks, indicating each model’s stability compared to the baseline. In general, different models demonstrate various capability in handing corruption data. Multi-attribute prediction VLM (InternVL2.5-2B) demonstrates superior stability compared to single-attribute CV models in most cases, especially when distinguishing building type and surface material and when encountering occlusion and brightness variations. CNN models’ stability performs comparably to more advanced models in the tasks of number of floors prediction, while ViT model performs superior in handling data with Gaussian noise. This outcome implies that additional domain-specific constraints or specialized training strategies might be required to enhance performance on crowdsourced image data.
Table 9: Relative prediction error under different corruptions and perturbations for different objective attributes.
Attribute Model Occlusion Motion Noise Brightness Relative mCE
Building type ResNet50 1.00 1.00 1.00 1.00 1.00
ResNet101 0.95 1.14 0.90 1.10 1.02
ViT16 1.09 1.22 0.59 1.60 1.13
InternVL2.5-2B 0.51 0.65 0.80 0.64 0.65
Surface material ResNet50 1.00 1.00 1.00 1.00 1.00
ResNet101 1.00 0.84 0.94 0.78 0.89
ViT16 1.61 1.01 0.79 1.64 1.26
InternVL2.5-2B 0.45 0.59 0.67 0.51 0.56
Number of floors ResNet50 1.00 1.00 1.00 1.00 1.00
ResNet101 1.05 0.91 0.82 0.89 0.92
ViT16 0.95 0.82 0.27 2.05 1.02
InternVL2.5-2B 1.07 1.44 0.86 0.85 1.05
Building age ResNet50 1.00 1.00 1.00 1.00 1.00
ResNet101 0.58 0.85 0.95 0.81 0.80
ViT16 0.40 0.62 0.70 1.99 0.93
InternVL2.5-2B 0.15 0.72 1.44 0.56 0.72
In conclusion, building on insights from prior CV techniques, VLMs not only demonstrate robust and generalizable features for tackling diverse tasks based on crowdsourced data, but they also represent a promising framework for large-scale or cross-regional implementations that demand multi-feature prediction and flexible adaptation. Moreover, incorporating more domain-specific and diverse, high-quality data can further enhance the framework’s performance and facilitate broader adoption in real-world scenarios.
4.4 Image labeling and captioning
Figure 13: Comparison of OSM building data (left) and the building attributes inferred using our method (right) in Washington D.C., illustrating attributes: building type, surface material, number of floors, and age. Data: (c) OpenStreetMap contributors.
Detected buildings across seven global cities, introduced in Section 4.1, are subsequently processed by the fine‐tuned VLM to generate objective attributes and captions. Overall, data for half a million buildings are enriched using 1.2 million images, each linked to its geographical location. For buildings with multiple observations, the most frequently assigned categories are retained. Figure 13 compares the availability of building properties before and after enrichment in Washington, D.C. The proposed approach effectively enhances building‐level information, particularly for surface material and building age, wherever street‐level imagery is available. Additionally, Table 10(d) presents the distribution of class labels for each attribute across the 1.2 million‐building dataset.
Beyond the predefined labels, our dataset includes text annotations for each building image, providing a richer source of information for categorizing architectural features. These captions capture intricate details beyond standard classifications, including facade styles, structural elements, and mixed-use characteristics, offering a more nuanced understanding of urban form. By extracting key descriptors, Figure 14 showcases examples of mixed-use buildings and diverse facade styles identified in Washington, D.C., and San Francisco. This methodology introduces additional dimensionalities for architectural feature analysis, allowing for more detailed characterizations of urban landscapes. Moreover, it facilitates fine-grained comparisons across cities, helping to reveal and interpret regional architectural trends and stylistic variations.
Figure 14: Spatial distribution of mixed-use buildings (top) and facade styles (bottom) in different cities. Data: (c) OpenStreetMap contributors.
5 Discussion
5.1 Application of building image dataset
Despite the centrality of objective building attributes in urban analytics, their scarcity still persists across cities (Biljecki et al., 2023). Our open framework OpenFACADES, addresses this gap by utilizing SVI — an urban sensing modality that captures pedestrian-scale visual information — together with building data to develop an open-sourced MLLM framework for unified attribute extraction and semantic description. The methodology begins by integrating crowdsourced SVI metadata with geometrical building data using isovist analysis to identify relevant images. Buildings are then detected based on their angles of view within image space, followed by an automated process of reprojecting and filtering them into individual building images. Lastly, a subset of this dataset is used to construct an image-text dataset designed for three tasks for VLM fine-tuning: single-word Q&A, multi-attribute Q&A, and captioning. Our experiments indicate that the fine-tuned VLM demonstrates strong performance in multi-attribute prediction, surpassing CV models and outperforming zero-shot ChatGPT-4o baselines. Deploying the VLM at scale, we annotate and release data of half a million buildings with both objective attributes and textual descriptions, derived from 1.2 million images across seven global cities, contributing to a scalable and automated approach for building property enrichment.
Our study directly features three main contributions to building research. First, our methodology detects holistic building facades and reprojects them into undistorted individual images, ensuring comprehensive visual coverage while reducing the uncertainty inherent in panoramic imagery. This pipeline can be integrated with existing methods to detect buildings from diverse viewing angles and associate them with geolocation, enabling nuanced and holistic observation for exterior modeling (Zhang et al., 2021), facade material segmentation (Tarkhan et al., 2025), and window-to-wall ratio calculation (De Simone et al., 2024). Second, this work introduces an inclusive and efficient pipeline to utilize both crowdsourced data and open-sourced LLMs for street-level research. This pipeline not only overcomes the challenge of relying on proprietary datasets, but also circumvents the high costs and limited adaptability associated with proprietary LLM APIs, making advanced analytical techniques more accessible and reproducible to the research community. Future studies might adjust the pipeline to customized tasks to incorporate fine-grained visual information with tailored building data based on their objectives, such as building conditions (Zou and Wang, 2021), human perceptual indicators (Liang et al., 2024) and seismic structural types (Pelizari et al., 2021).
Third, we present unified benchmark VLMs that perform multi‐task learning on building facades, generating descriptive captions while maintaining robust multi‐class predictions of objective attributes. In particular, we:
  • Explore the capabilities of VLMs through zero‐shot settings, varying scales of training data, and applications across different cities and attribute categories. Our findings show that fine‐tuning is essential to enhance performance in current VLMs, especially given the inherent quality issues in crowdsourced building data. Augmenting computer‐generated labels with ground‐truth annotations significantly improves model performance. Furthermore, our experiments demonstrate that efficient training can be achieved by using only 5% of the available crowdsourced data (approximately 1,000 image–text pairs), or by supplementing sufficient ground truth building data with minimal GPT-generated content, while still yielding robust performance gains of multi-attribute prediction.
  • Apply the method at scale by generating labels and captions for half a million buildings in eight cities, laying a foundation for future urban analyses. For instance, integrating these labeled data with geospatial information can add new dimensions to urban functional zone classification (Zhang et al., 2023), including potential insights into 3D functional zoning (Lin et al., 2024). The unified model also infers multi‐dimensional building properties relevant for applications such as modeling building electricity consumption (Rosenfelder et al., 2021), estimating material stocks (Raghu et al., 2023), and assessing structural risk (Wang et al., 2021). Additionally, captions offer an extra layer of information about building facades, enabling the identification of mixed‐use buildings or stylistic variations. This linguistic data holds promise for exploring urban identity, supporting text‐image‐based generative design, and serving as an additional feature layer in multimodal model training.
  • Demonstrate enhanced generalizability and robustness by comparing VLMs with CV models. Owing to larger parameter counts and extensive pretraining on general‐purpose data, VLMs excel in predicting objective building attributes for both unseen data from culturally different region and images with variable conditions. This result provides a convincing avenue for future cross‐city analyses and crowdsourced data research, where broad adaptation and resilience to heterogeneous imagery remain critical challenges.
5.2 Limitations and future works
Despite the advancements presented in this study, limitations remain. First, while this study incorporates captioning data for fine-tuning VLMs, these captions are generated using commercial state-of-the-art LLMs rather than human-labeled ground truth, leaving their accuracy and reliability unverified. A systematic human evaluation would be valuable for future research to assess captioning quality, consistency, and semantic accuracy. Additionally, leveraging open-access models offers a more sustainable approach for scalable dataset expansion in future studies. Knowledge distillation — where a smaller model learns from a larger teacher model — presents a promising solution for self-supervised learning, enabling broader generalization across diverse urban settings and more efficient adaptation to building-related tasks.
Second, while the fine-tuned model exhibits strong generalizability across cities, the quality of crowdsourced data remains a crucial factor (Biljecki et al., 2023; Hou and Biljecki, 2022). Although this study incorporates multiple strategies to mitigate data quality issues — such as isovist-based vantage point selection, building data harmonization, and feature-based image filtering — challenges persist. These include incorrect or incomplete building labels, inconsistent geometry information, non-standardized image format, and misaligned image coordinates, all of which contribute to different sources of uncertainty. Future work should focus on enhancing dataset reliability through improved data filtering mechanisms. Automated repetition detection, heuristic rule-based filtering, and uncertainty-aware sampling could refine image selection and minimize inconsistencies in building attribute annotations (Chen et al., 2024b).
6 Conclusion
This comprehensive study advances spatial data infrastructures and urban data science by introducing a novel framework, OpenFACADES, which leverages volunteered geographic information to enrich building profiles on a global scale using street-level imagery and multimodal large language models. We harvest multimodal crowdsourced data and apply isovist analysis, object detection, and a tailored reprojection method to geolocate and acquire holistic building images, thereby establishing a comprehensive global building image dataset. A selection of this open dataset is then utilized for fine-tuning large vision-language models (VLMs), enabling large-scale enrichment of building profiles through multi-attribute prediction (e.g., material, function) and open-vocabulary captioning.
Our findings demonstrate that VLMs outperform conventional CNN-based models and zero-shot GPT-4o baselines in predicting building attributes while generating linguistically grounded descriptions. This methodological advancement has enabled the creation of a large-scale dataset covering half a million buildings across seven global cities, addressing critical gaps in building data availability for urban analytics. By bridging the limitations of existing datasets, this framework provides a scalable solution for capturing multi-dimensional fine-grained architectural details and urban morphological characteristics. The enriched dataset further facilitates a more nuanced and expansive exploration of urban environments, with potential applications in energy modeling, risk assessment, and sustainable development.
Beyond its immediate applications, we envision this framework as a foundation for comprehensive building profiling, capturing not only physical attributes but also the socio-economic and cultural narratives embedded within the built environment. This advancement has significant implications for urban research, including large-scale built environment analysis, building simulation, and policy-driven planning strategies.
Acknowledgments
This research is part of the project Large-scale 3D Geospatial Data for Urban Analytics, which is supported by the National University of Singapore under the Start Up Grant. This research is part of the project Multi-scale Digital Twins for the Urban Environment: From Heartbeats to Cities, which is supported by the Singapore Ministry of Education Academic Research Fund Tier 1. The first author acknowledges the NUS Graduate Research Scholarship granted by the National University of Singapore (NUS). We thank the members of the NUS Urban Analytics Lab for the discussions. We also acknowledge the contributors of OpenStreetMap, Mapillary and other platforms for providing valuable open data resources and code that support street-level imagery research and applications.
Appendix A Building dataset supplementary
Table 10: Detailed breakdown of building attributes in the annotated image dataset using the OpenFACADES framework.
(a) Building type
apartments house garage industrial religious retail hotel office public education
Amsterdam 135,233 133,794 2,127 10,386 937 18,074 2,703 3,281 2,329 2,858
Helsinki 8,165 5,371 678 983 101 995 401 1,227 1,393 749
Berlin 113,763 117,604 3,195 6,315 1,438 10,635 7,205 5,925 13,396 4,392
San Francisco 31,240 42,645 315 1,241 538 11,151 832 2,171 647 677
Houston 14,553 177,931 13,223 13,568 1,916 10,395 469 2,019 2,788 1,438
Washington D.C. 20,389 158,230 540 1,435 1,772 6,935 1,005 2,703 1,667 3,004
Manila 13,989 11,915 110 1,892 448 15,257 281 1,226 1,314 1,890
All 337,332 647,490 20,188 35,820 7,150 73,442 12,896 18,552 23,534 15,008
(b) Surface material
brick wood plaster concrete metal glass stone
Amsterdam 95,241 2,606 6,960 2,081 2,802 1,278 460
Helsinki 2,225 1,571 2,910 894 419 445 134
Berlin 15,708 3,697 104,438 5,992 2,227 2,160 1,441
San Francisco 2,968 9,324 19,823 864 333 654 447
Houston 24,740 45,756 13,359 2,554 3,327 555 819
Washington D.C. 58,160 13,063 10,530 823 218 840 1,349
Manila 467 1,915 7,638 11,503 1,002 782 205
All 199,509 77,932 165,658 24,711 10,328 6,714 4,855
(c) Number of floors
1 2 3 4-5 6-7 8-9 10-12 13-15 16-20 above 20
Amsterdam 6,939 38,133 19,209 45,239 1,142 227 320 150 39 12
Helsinki 2,281 2,076 740 1,951 1,289 118 98 19 7 10
Berlin 26,506 47,651 12,863 37,768 8,485 593 1,299 143 249 56
San Francisco 1,583 13,216 14,598 3,647 581 199 243 37 121 175
Houston 64,350 20,896 4,720 585 149 70 134 47 64 94
Washington D.C. 6,951 61,847 11,905 2,518 590 390 669 45 9 10
Manila 2,190 9,726 5,330 4,274 929 104 426 81 138 303
All 110,800 193,545 69,365 95,982 13,165 1,701 3,189 522 627 660
(d) Building age
Before 1800 1800-1900 1900-1950 1950-1975 1975-2000 After 2000
Amsterdam 847 20,853 19,203 34,470 27,067 2,402
Helsinki 16 707 818 3,755 2,684 617
Berlin 169 17,433 19,150 65,635 30,221 3,029
San Francisco 5 9,003 13,488 5,406 6,128 371
Houston 0 1,980 7,048 39,554 41,376 1,152
Washington D.C. 80 22,068 33,950 20,307 7,976 582
Manila 0 110 278 5,778 17,031 317
All 1,117 72,154 93,935 174,905 132,483 8,470